OpenStudy (johnjules2000):

4m-5(m+2)=1 Will somebody please help me solve this, solve for m and the 5 is subtraction.

OpenStudy (johnjules2000):

thank you

OpenStudy (anonymous):

isn't it m=-11

OpenStudy (the_fizicx99):

Typo, 4m-5(m+2)=1 4m - 5m -10 = 1 -m - 10 = 1 +10 +10 -m = 11 /-1 /-1 M= -11 4(-11)-5(-11+2)=1 -44 +55 + 10 = 1 11 - 10 = 1 1 = 1 so this is true

OpenStudy (johnjules2000):

how did you figure that out? i might need to know for future reference.

OpenStudy (the_fizicx99):

I applied distributive property to -5(m+2) 4m - 5m -10 = 1 Then you combine like terms, 4m-5m = -1m or just -m -m - 10 = 1 You want to isolate m, so the inverse of subtraction is addition, add 10 to both sides -m = 11 m can't be negative so divide by -1 -m = 11 /-1 /-1 m = -11